Calculate the pH of an aqueous solution of sodium hydrogen phosphite, NaHPO3.

For phosphorous acid, H2PO3, pKa'= 2.00 and pLaTeX: K_a^{''}Ka″ = 6.59.
Answer #1

Dissociation of H2PO3 in water takes place in two steps.

1. H2PO3 (aq) + H2O (l)  \rightleftharpoons H3O + (aq) + HPO3- (aq) pKa ' = 2.00

2. HPO3- (aq) + H2O (l)  \rightleftharpoons H3O + (aq) + PO32- (aq) pKa " = 6.59

From above reactions, it is clear that HPO3- is an intermediate form of H2PO3 and PO32-.

We know that, pH of intermediate form of diprotic acid is calculated by using formula pH = 1/2 ( pK 1 + pK 2)

\therefore pH of aqueous solution of  NaHPO3 = 1/2 (  pKa ' + pKa " )

= 1/2 ( 2.00 + 6.59 )

= 1/2 ( 8.59 )

= 4.295

= 4.30

ANSWER : pH of aqueous solution of  NaHPO3 =4.30